# Why Transformers Need Adam: A Hessian Perspective

### v1

@@ 3.3 Initial Theoretical Results
Complexity $\tilde{O}(\kappa)$ $\tilde{O}(\max_{l\in[L]}\kappa_{l})$ $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$
Table 2: The complexity of GD, BGD and Adam for minimizing a strongly convex quadratic function with block diagonal Hessian. The symbol ✗ means non-convergence. $\kappa$ and $\kappa_{l}$ denote the condition number of the full Hessian and the block submatrix, respectively. $r:=\max_{l\in[L]}\left(\frac{C_{l,2}}{C_{l,1}}\right)^{2}$ and $C_{l,1}$, $C_{l,2}$ are constants in Assumption 1.
@@
It is valid to ask whether the complexity upper bound $\kappa_{Adam,l}=r\kappa_{l}$ can be tightened, e.g., remove the factor of $r$. We point out it would be difficult if there is no extra structure on $H_{l}$. A key technical step is to bound the condition number of the preconditioned matrix $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$. Intuitively, a diagonal preconditioner of $H_{l}$ is expected to be powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant. Unfortunately, it is unclear whether these structures hold in real Transformers. Without any assumption on $H_{l}$, we find that the diagonal preconditioner of $D_{Adam}^{0}$ could increase the condition number. For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $3.9\kappa_{1}$, $14.1\kappa_{2}$, $4.2\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials). It would be interesting to explore if there are special structures of $H_{l}$ in real Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it. We leave it as a future direction.
@@
4 Related Works
On The failure of SGD on Transformers
There is extensive literature studying the difficulties of Transformer training. We summarize these works in Appendix A. We here focus on the literature that explores why SGD fails on Transformers. One representative hypothesis is that SGD fails because it cannot handle the heavy-tailed stochastic noise in language tasks (Zhang et al., 2020). However, Chen et al. (2021); Kunstner et al. (2023) reported that the gap between Adam and SGD maintains even in the full-batch case with no stochasticity, so there might be other reasons. Further, SGD performs poorly on Vision Transformers on ImageNet (Figure LABEL:fig:nlp_figure), so the data modality (e.g., language or image tasks) might not be as crucial as the architecture. Zhang et al. (2019c) showed that Transformers have the “unbounded smoothness" issue and SGD with gradient clipping performs better than SGD in this case. Although clipping is an effective trick, it does not save SGD from failure as we still observe a huge gap between clipped SGD and Adam  [^3] For all NLP tasks, clipping is performed after backpropagation. So in Figure LABEL:fig:nlp_figure, SGD in NLP tasks essentially refers to clipped SGD. . So there might be other reasons that hamper SGD. Different from these works, we find SGD fails because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
Understanding of Adam.
There was once a long-standing debate on the possible divergence of Adam (Reddi et al., 2018). The convergence for the unmodified versions is later established in (Shi et al., 2020; Zhang et al., 2022b) for RMSprop and Adam. More convergence analyses of general adaptive gradient methods are listed in Appendix A. We here focus on the literature that explores the benefit of Adam. Xie et al. (2022) show that Adam can help avoid saddle points, which is an orthogonal direction to this work. Wang et al. (2022a); Crawshaw et al. (2022); Li et al. (2023) show that Adam and its variant outperform SGD under relaxed smoothness conditions, based on the intuition that Adam can adaptively change its learning rate along iteration (over time). We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes fails while Adam works well (Figure LABEL:fig_quadratic). This indicates that the benefit of Adam is not merely due to its ability to adaptively change the learning rate over time, and there are other reasons for Adam’s success. We show that an important benefit of Adam is its ability to handle the heterogeneity across blocks (over space).
Bernstein et al. (2018); Wu et al. (2020); Kunstner et al. (2023); Liu et al. (2023) build a relation between Adam and the sign-based methods. Wu et al. (2020) further showed that sign-based methods can be effective when the Hessian is diagonal and satisfies several other properties. However, as put by the authors, it seems “unclear to what extent these properties hold for real problems". Pan and Li (2023) numerically found that the Adam can reduce the directional sharpness along trajectories, while its relation to fast convergence remains mysterious. A recent work (Jiang et al., 2023) point out that Adam biases the trajectories towards regions where Hessian has “uniform diagonal entries" while SGD cannot. The distribution of Hessian diagonal entries is also investigated in (Liu et al., 2023). The theory in (Jiang et al., 2023) implies that Adam is faster when the Hessian is diagonal. However, as argued above, it is unclear whether the diagonal Hessian commonly holds in real problems. In fact, we find the Hessian is closer to a block-diagonal (instead of pure diagonal) structure on some small Transformers. In these cases, blockwise eigenvalues carry more information than diagonal entries, providing extra details such as the location of eigenvalues. We find that these extra details are important for distinguishing Adam and SGD.
Hessian Spectrum Analysis.
There are several important attempts to explore the Hessian spectrum of MLPs and CNNs. Sagun et al. (2016, 2017); Chaudhari et al. (2019) found that the Hessian spectra of MLPs and CNNs consist of a “bulk" together with a few “outliers". Papyan (2020); Wu et al. (2020); Liao and Mahoney (2021) further characterized the bulks and outliers in theory. Papyan (2018, 2019) numerically built the relation between these "outliers" and the Gauss-Newton matrix. Sankar et al. (2021) numerically explored the relation between Hessian of CNNs and Gauss-Newton matrix in each layer. They further showed that most CNN layers contribute similarly to the overall loss surface. We find that this result is restricted to CNNs and does not hold on Transformers due to the heterogeneity. Gur-Ari et al. (2018) showed that for MLPs and CNNs, gradient descent converges to a small subspace spanned by a few top eigenvectors of the Hessian. Yao et al. (2018); Zhang et al. (2019b) explored the relation between the Hessian spectrum of CNNs and some training phenomena such as the effect of batch sizes. Ghorbani et al. (2019); Yao et al. (2020) focused on explaining the effectiveness of techniques such as BatchNorm. Note that all these works are restricted to MLPs and CNNs, while we study the Hessian of Transformers (in addition to CNNs and MLPs) as well as its impacts on different optimizers.
5 Conclusion
In this work, we investigate the failure mode of SGD on Transformers through the lens of Hessian. By numerically explore various Transformers, CNNs, MLPs and quadratic problems, we establish a phenomenon called block heterogeneity in Hessian and link it to the failure of SGD. We point out that SGD fails because it applies one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. Initial theory are provided to support the claim.
Appendix A More Related Works
On the difficulties of Transformer training.
Transformers are known to be difficult to train. Researchers have attributed the training difficulties to various phenomena in different components of Transformers, including: the logits divergence or the rank degeneracy in the outputs of attention layers [Dong et al., 2021, Noci et al., 2022, Wortsman et al., 2023, Zhai et al., 2023, Dehghani et al., 2023, Chowdhery et al., 2023]; the growth of parameter norm in attention layers [Merrill et al., 2020]; over-reliance on residue branches [Liu et al., 2020]; and some negative impact of layer norm [Chen et al., 2018, Zhang et al., 2019a, Huang et al., 2020]. These phenomena have a strong correlation with gradient vanishing or explosion in Transformers [Zhang et al., 2019a, Liu et al., 2020, Huang et al., 2020, Xiong et al., 2020, Noci et al., 2022, Wang et al., 2022b, Wortsman et al., 2023, Molybog et al., 2023], which leads to training difficulties.
Several solutions have been proposed. Liu et al. [2020] numerically observed that adaptive gradient methods can (partly) overcome gradient vanishing by giving “consistent update magnitude", while it seems unclear how consistent update magnitude would help optimization in principle. Researchers further develop training tricks such as warmup learning rate [Liu et al., 2019, Xiong et al., 2020], temperature scaling [Noci et al., 2022], better initialization [Zhang et al., 2019a, Huang et al., 2020, Wang et al., 2022b, Bachlechner et al., 2021, Yang et al., 2022], and variants of Layer Norm [Nguyen and Salazar, 2019, Wang et al., 2019, Xiong et al., 2020, Wang et al., 2022b, Dehghani et al., 2023]. Recent researchers also suggest using z-loss regularization [Chowdhery et al., 2023, Yang et al., 2023] and tuning hyperparameters of Adam [Zhang et al., 2022b, Wortsman et al., 2023]. All these tricks can help mitigate gradient explosion or vanishing. Nevertheless, training large-scale Transformers remains challenging [Zhang et al., 2022a, Zeng et al., 2022, Wortsman et al., 2023, Molybog et al., 2023, Chowdhery et al., 2023]. Different from all aforementioned works, we investigate the training difficulties of Transformers through the eigenvalues of Hessian. We establish a strong correlation between the blockwise Hessian spectra of Transformers and the failure of SGD. We realize that our attempt is just a first step towards understanding Transformer training, and we believe there is rich information hidden in Hessian and we leave more fine-grained analysis as future works.
Convergence analysis of adaptive gradient methods
There is extensive convergence analysis for adaptive gradient methods. For instance, researchers study the convergence of AMSGrad [Reddi et al., 2018, Zhou et al., 2018], RMSprop [Zaheer et al., 2018], AdaFom [Chen et al., 2019], AdaBound [Luo et al., 2018], and Adam with iterate-dependent hyperparameters [Zou et al., 2019, Chen et al., 2022, Gadat and Gavra, 2022]. The convergence of Adam is also explored in [Défossez et al., 2022, Wang et al., 2023a]. There is also an active line of theoretical research on the convergence of AdaGrad [Duchi et al., 2011], we recommend [Wang et al., 2023b] for more detailed introduction. In this work, we do not focus on the convergence analysis. Rather, we explore the quantitative difference between the loss landscape of CNNs and Transformers and how it impact the behaviors of SGD and Adam.
Appendix B More Preliminaries
B.1 Preliminaries on Optimizers
@@
B.2 Preliminaries on the Stochastic Lanczos Quadrature Method
@@
Approximation of the Hessian spectrum can be formulated as a trace estimation problem, as introduced in [Lin et al., 2016, Ubaru et al., 2017]. First, the spectrum (eigenvalue density) of Hessian $H$ can written as: $\phi(t)=\frac{1}{d}\sum_{i=1}^{d}\delta\left(t-\lambda_{i}\right)$, where $\lambda_{i}$ are the eigenvalues of $H$ and $\delta$ is the Dirac $\delta$-function. Then, we replace the delta functions by a Gaussian blurring function: $\phi(t)\approx g(t):=\frac{1}{d}\sum_{i=1}^{d}f\left(\lambda_{i}\right)$, where $f(\lambda):=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(t-\lambda)^{2}}{2\sigma^{2}}\right)$. By definition of matrix function, it is easy to see that $g(t)=\frac{1}{d}tr(f(H))$. As such, spectrum approximation could be formulated as a trace estimation problem, i.e., estimating $\frac{1}{d}\operatorname{tr}(f(H))$, where $H\in\mathbb{R}^{d\times d}$ is a real symmetric matrix.
Trace estimation problems could be solved efficiently by the Stochastic Lanczos Quadrature Method (SLQ) [Golub and Strakoš, 1994]. For the ease of readers, we provide a detailed description of SLQ in our context. SLQ consists of the following steps.
@@
We Approximate the trace of matrix function as $\frac{1}{d}tr(f(H))=\mathbb{E}(v^{T}f(H)v)\approx\frac{1}{n_{v}}\sum_{i}^{n_{v}}v_{i}^{T}f(H)v_{i}$, where $v=u/\|u\|_{2}$ and $u$ is a Rademacher random vector (each entry of $u$ independently takes ±$\pm 1$ with probability $1/2$). This step is called Huchinson’s estimation [Hutchinson, 1989].
Note that we can also replace the Rademacher random vector $u$ by a unit Gaussian vector (i.e., $u\sim N(0,I_{d\times d})$) and the unbiasedness still holds [Avron and Toledo, 2011]. In our implementation, we sample $u\sim N(0,I_{d\times d})$ because there is an efficient built-in PyTorch function for generating Gaussian vectors.
@@
We rewrite $v^{T}f(H)v$ as a Riemann-Stieltjes integral [Golub and Meurant, 2009]:
@@
We hope to design $\{(c_{j},x_{j})\}_{j=1}^{m}$ with a reasonable number of $m$ such that the estimation error is small. Fortunately, the Gaussian Quadrature method provides a generic design principle of $\{(c_{j},x_{j})\}_{j=1}^{m}$ [Golub and Meurant, 2009, Epperson, 2013]. It is proved that: when $f(\lambda)$ is not "too complicated" (e.g. $f(\lambda)$ is a polynomial), then there exists $\{(c_{j},x_{j})\}_{j=1}^{m}$ which gives a high quality estimation of integral (7). The required number of $m$ is related to "how complicated the $f(\lambda)$ is". Such $\{(c_{j},x_{j})\}_{j=1}^{m}$ are called the Gaussian Quadrature rules. $c_{j}$ and $x_{j}$ are called the "weights" and the "nodes" of the Gaussian Quadrature rules. A representative theorem is as follows: when $f(\lambda)$ is a polynomial with degree $<2m$, then the Gaussian Quadrature rules give the exact approximation of integral (7).
@@
Theorem 3 shows the existence of good $\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ and their general form. In fact, it is also shown that Gaussian Quadrature is optimal: no other$\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ can achieve zero approximation error for higher degree polynomials $f(\lambda)$ [Golub and Meurant, 2009]. However, it is often difficult to find these quadrature rules [Golub and Welsch, 1969]. There are at least three questions in sequel:
@@
Now suppose that we have found the orthogonal polynomials $\left\{p_{k}(x)\right\}_{k=1}^{m}$ w.r.t. $\mu$. Recall that any orthogonal polynomial has the following "three-term" recursion [Golub and Meurant, 2009]:
@@
[Golub and Meurant, 2009] For the Gaussian Quadrature, $\{x_{j}\}_{j=1}^{m}$ are the eigenvalues of $J_{m}$ and $\{c_{j}\}_{j=1}^{m}$ are the squares of the first elements of the normalized eigenvectors of $J_{m}$.
The proof of Theorem 4 is based on Christoffel-Darboux relation [Brezinski, 1990]. Now, the remaining question is: how to find the Jacobian matrix $J_{m}$ of a sequence of orthogonal polynomials w.r.t. an unknown measure $\mu$? Note that we no longer need to answer question 1) if $J_{m}$ is found, since $J_{m}$ is sufficient for us to find the Gaussian quadrature rules. However, it seems impossible to find $J_{m}$ if no information of $\mu$ is provided. The good news is: when the $\mu$ is specified as in (8), there exists an efficient way to find $J_{m}$.
@@
When $\mu$ is specified as in (8), $J_{m}$ can be exactly found in $m$ steps using the Lanczos algorithm [Lanczos, 1950], as shown in Algorithm 4. This method takes a real symmetric matrix as input and returns a tridiagonal matrix. It was originally proposed to solve eigenvalue problems. Later, researchers found a deep connection between the Lanczos algorithm and orthogonal polynomials, which further connects this method to the Gaussian quadrature. The method (of finding the Gaussian quadrature by the Lanczos algorithm) is called the Lanczos quadrature [Golub and Strakoš, 1994, Bai and Golub, 1996, Golub and Meurant, 2009]. An extremely elegant but highly nontrivial result is as follows:
@@
[Golub and Meurant, 2009] Given a real symmetric matrix $H\in\mathbb{R}^{d\times d}$ and an arbitrary vector $v\in\mathbb{R}^{d}$ with unit Euclidean norm, we define the measure $\mu$ as in (8) based on this $H$ and $v$. Then $m$ steps of the Lanzcos algorithm return the Jacobian matrix $J_{m}$ of orthogonal polynomials w.r.t. to $\mu$.
@@
The full procedure of SLQ is shown in Algorithm 5. We note that SLQ is efficient in theory. Ubaru et al. [2017] show that SLQ converges faster than any other polynomial expansion method for spectrum estimation (e.g., Chebyshev methods used in [Adams et al., 2018]). See [Ubaru et al., 2017, Theorem 4.1] for a formal statement.
We remark that there are at least four versions of the Lanczos algorithm in Step 3. Here, we adopt the version in Algorithm 4 since it is known to be the most numerically stable version [Cullum and Willoughby, 2002, Saad, 2011, Wikipedia, 2023]. Throughout this work, we choose $f\left(\cdot\right)$ as the Gaussian blurring function $f(\lambda):=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(t-\lambda)^{2}}{2\sigma^{2}}\right)$ for spectrum approximation. We plot the spectrum by sweeping $t$ from the minimal node to the maximal node in Gaussian Quadrature rules.
@@
Appendix C More Eperimental Details
C.1 Implementation Details on SLQ

### v2

@@ 3.2 Initial Theoretical Results
Complexity $\tilde{O}(\kappa)$ $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$
4 Conclusion
In this work, we explore why SGD performs badly on Transformers. By numerically exploring various neural nets and quadratic problems, we establish a phenomenon called block heterogeneity in Hessian and link it to the deficiency of SGD. Initial theory is provided to support the claim.
Appendix A Related Works
On the deficiency of SGD on Transformers
There is an active line of works that explores why SGD performs badly on Transformers. One representative hypothesis is that SGD cannot handle the heavy-tailed stochastic noise in language tasks [106]. However, Chen et al. [14], Kunstner et al. [46] reported that the gap between Adam and SGD maintains even in the full-batch case with no stochasticity, so there might be other reasons. Further, SGD performs poorly on Vision Transformers on ImageNet (See Figure LABEL:fig:nlp_figure. Also see [93] for more evidence), so the data modality (e.g., language or vision tasks) might not be as crucial as the architecture. [105] showed that NLP tasks have “unbounded smoothness" issue and SGD with gradient clipping performs better than SGD in this case. Although clipping is an effective trick, we still observe a huge gap between clipped SGD and Adam  [^4] For all NLP tasks, clipping is performed immediately after backpropagation. So in Figure LABEL:fig:nlp_figure, SGD in NLP tasks essentially refers to clipped SGD. . So there might be other reasons that hamper SGD. Different from these works, we find SGD performs badly because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
Understanding of Adam.
There was once a long-standing debate on the possible divergence of Adam [72]. The convergence for the unmodified versions is later established in [78, 108] for RMSprop and Adam. More convergence analyses of general adaptive gradient methods are listed later in this section. We here focus on the literature that explores the benefit of Adam. Xie et al. [94] show that Adam can help avoid saddle points, which is an orthogonal direction to this work. Wang et al. [85], Crawshaw et al. [19], Li et al. [50] show that Adam and its variant outperform SGD under relaxed smoothness conditions, based on the intuition that Adam can adaptively change its learning rate along iteration (over time). We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes performs badly while Adam works well (Figure LABEL:fig_quadratic). This indicates that the benefit of Adam is not merely due to its ability to adaptively change the learning rate (over time), and there are other reasons for Adam’s success. We show that an important benefit of Adam is its ability to handle the heterogeneity across blocks (over space).
Recent works [9, 92, 46, 53, 3] build a relation between Adam and the sign-based methods. Wu et al. [92] further showed that sign-based methods can be effective when the Hessian is diagonal and satisfies several other properties. However, as put by the authors, it seems “unclear to what extent these properties hold for real problems". Pan and Li [66] numerically found that the Adam can reduce the directional sharpness along trajectories, while its relation to fast convergence remains mysterious. A recent work [43] point out that Adam biases the trajectories towards regions where Hessian has “uniform diagonal entries" while SGD cannot. The distribution of Hessian diagonal entries is also investigated in [53]. The theory in [43] implies that Adam is faster when the Hessian is diagonal. However, as argued above, it is unclear whether the diagonal Hessian structure commonly holds in real problems. In fact, we find the Hessian is closer to a block-diagonal (instead of pure diagonal) structure on some small Transformers. In these cases, blockwise eigenvalues carry more information than diagonal entries, providing extra details such as the location of eigenvalues. We find that these extra details are important for distinguishing Adam and SGD.
Hessian Spectrum Analysis.
There are several important attempts to explore the Hessian spectrum of MLPs and CNNs. Early works [75, 76, 12] found that the Hessian spectra of MLPs and CNNs consist of a “bulk" together with a few “outliers". Papyan [69], Wu et al. [92], Liao and Mahoney [51] further characterized the bulks and outliers in theory. Papyan [67, 68] numerically built the relation between these "outliers" and the Gauss-Newton matrix. Sankar et al. [77] numerically explored the relation between Hessian of CNNs and Gauss-Newton matrix in each layer. They further found that most CNN layers contribute similarly to the overall loss surface. We find that this result is restricted to CNNs and does not hold on Transformers due to the heterogeneity. Gur-Ari et al. [39] showed that for MLPs and CNNs, gradient descent converges to a small subspace spanned by a few top eigenvectors of the Hessian. Yao et al. [98], Zhang et al. [104] explored the relation between the Hessian spectrum of CNNs and some training phenomena such as the effect of batch sizes. Ghorbani et al. [32], Yao et al. [99] focused on explaining the effectiveness of techniques such as BatchNorm. Note that all these works are restricted to MLPs and CNNs, while we study the Hessian of Transformers (in addition to CNNs and MLPs) as well as its impacts on different optimizers.
On the difficulties of Transformer training.
Transformers are known to be difficult to train. Researchers have attributed the training difficulties to various phenomena in different components of Transformers, including: the logits divergence or the rank degeneracy in the outputs of attention layers [26, 65, 91, 102, 24, 17]; the growth of parameter norm in attention layers [61]; over-reliance on residue branches [55]; and some negative impact of layer norm [15, 103, 41]. These phenomena have a strong correlation with gradient vanishing or explosion in Transformers [103, 55, 41, 95, 65, 88, 91, 62], which leads to training difficulties.
Several solutions have been proposed. Liu et al. [55] numerically observed that adaptive gradient methods can (partly) overcome gradient vanishing by giving “consistent update magnitude", while it seems unclear how consistent update magnitude would help optimization in principle. Researchers further develop training tricks such as warmup learning rate [54, 95], temperature scaling [65], better initialization [103, 41, 88, 5, 97], and variants of Layer Norm [64, 89, 95, 88, 24]. Recent researchers also suggest using z-loss regularization [17, 96] and tuning hyperparameters of Adam [108, 91]. All these tricks can help mitigate gradient explosion or vanishing. Nevertheless, training large-scale Transformers remains challenging [107, 101, 91, 62, 17]. Different from all aforementioned works, we investigate the training difficulties of Transformers through the eigenvalues of Hessian. We establish a strong correlation between the blockwise Hessian spectra of Transformers and the poor performance of SGD. We realize that our attempt is just a first step towards understanding Transformer training, and we believe there is rich information hidden in Hessian and we leave more fine-grained analysis as future works.
Convergence analysis of general adaptive gradient methods
There is extensive convergence analysis for adaptive gradient methods. For instance, researchers study the convergence of AMSGrad [72, 109], RMSprop [100], AdaFom [16], AdaBound [57], and Adam with iterate-dependent hyperparameters [110, 13, 31]. The convergence of Adam is also explored in [23, 86]. There is also an active line of theoretical research on the convergence of AdaGrad [28], we recommend [87] for more detailed introduction. In this work, we do not focus on the convergence analysis. Rather, we explore the quantitative difference between the loss landscape of CNNs and Transformers and how it impact the behaviors of SGD and Adam.
Appendix B More Results and Discussions
Performance comparison of AdamW and SGD on different Architectures.
Here, we show the performance comparison of AdamW and SGD on different models. All the vision models are trained on ImageNet. Language models are trained on different English corpus. See Appendix D.1 for more implementation details.
Block-diagonal structure in the existing literature.
We remark that Collobert [18], Roux et al. [73], Martens and Grosse [60] also observed the block-diagonal structure in (approximated) Hessian of small-scaled MLPs. We restate their findings in Figure LABEL:fig_block_diagnal_existing. [18, Section 7] further theoretically proved that the block diagonal structure stems from (i) the layer-by-layer structure and (ii) the Cross-Entropy loss. These results suggest that block-diagonal structure might be common in NNs.
Blockwise spectra for quadratic models in Section 3.1.
We here visualize the blockwise spectrum for the quadratic models in Case 1 and Case 2. These spectra are collected from GPT2 and ResNet18, respectively.
@@ Appendix B More Results and Discussions
It is valid to ask whether the complexity upper bound in Theorem 1 $\kappa_{Adam,l}=r\kappa_{l}$ can be tightened, e.g., improve the factor of $r$. We point out it would be difficult if there is no extra structure on $H_{l}$. A key technical step is to bound the condition number of the preconditioned matrix $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$. Intuitively, a diagonal preconditioner of $H_{l}$ is powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant [30]. Unfortunately, it is unclear whether these structures hold in Transformers. Without any assumption on $H_{l}$, we find that the diagonal preconditioner of $D_{Adam}^{0}$ could increase the condition number. For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $7.09\kappa_{1}$, $18.98\kappa_{2}$, $18.76\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials). It would be interesting to explore if there are special structures of $H_{l}$ in Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it. We leave it as a future direction.
@@
Appendix C More Preliminaries
C.1 Preliminaries on Optimizers
@@
C.2 Preliminaries on the Stochastic Lanczos Quadrature Method
@@
Approximation of the Hessian spectrum can be formulated as a trace estimation problem, as introduced in [52, 83]. First, the spectrum (eigenvalue density) of Hessian $H$ can written as: $\phi(t)=\frac{1}{d}\sum_{i=1}^{d}\delta\left(t-\lambda_{i}\right)$, where $\lambda_{i}$ are the eigenvalues of $H$ and $\delta$ is the Dirac $\delta$-function. Then, we replace the delta functions by a Gaussian blurring function: $\phi(t)\approx g(t):=\frac{1}{d}\sum_{i=1}^{d}f\left(\lambda_{i}\right)$, where $f(\lambda):=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(t-\lambda)^{2}}{2\sigma^{2}}\right)$. By definition of matrix function, it is easy to see that $g(t)=\frac{1}{d}tr(f(H))$. As such, spectrum approximation could be formulated as a trace estimation problem, i.e., estimating $\frac{1}{d}\operatorname{tr}(f(H))$, where $H\in\mathbb{R}^{d\times d}$ is a real symmetric matrix.
Trace estimation problems could be solved efficiently by the Stochastic Lanczos Quadrature Method (SLQ) [35]. For the ease of readers, we re-organize and summarize the existing literature ([35, 83, 32]) and provide a detailed description of SLQ in our context. SLQ consists of the following steps.
@@
We Approximate the trace of matrix function as $\frac{1}{d}tr(f(H))=\mathbb{E}(v^{T}f(H)v)\approx\frac{1}{n_{v}}\sum_{i}^{n_{v}}v_{i}^{T}f(H)v_{i}$, where $v=u/\|u\|_{2}$ and $u$ is a Rademacher random vector (each entry of $u$ independently takes ±$\pm 1$ with probability $1/2$). This step is called Huchinson’s estimation [42].
Note that we can also replace the Rademacher random vector $u$ by a unit Gaussian vector (i.e., $u\sim N(0,I_{d\times d})$) and the unbiasedness still holds [4]. In our implementation, we sample $u\sim N(0,I_{d\times d})$ because there is an efficient built-in PyTorch function for generating Gaussian vectors.
@@
We rewrite $v^{T}f(H)v$ as a Riemann-Stieltjes integral [34]:
@@
We hope to design $\{(c_{j},x_{j})\}_{j=1}^{m}$ with a reasonable number of $m$ such that the estimation error is small. Fortunately, the Gaussian Quadrature method provides a generic design principle of $\{(c_{j},x_{j})\}_{j=1}^{m}$ [34, 29]. It is proved that: when $f(\lambda)$ is not "too complicated" (e.g. $f(\lambda)$ is a polynomial), then there exists $\{(c_{j},x_{j})\}_{j=1}^{m}$ which gives a high quality estimation of integral (7). The required number of $m$ is related to "how complicated the $f(\lambda)$ is". Such $\{(c_{j},x_{j})\}_{j=1}^{m}$ are called the Gaussian Quadrature rules. $c_{j}$ and $x_{j}$ are called the "weights" and the "nodes" of the Gaussian Quadrature rules. A representative theorem is as follows: when $f(\lambda)$ is a polynomial with degree $<2m$, then the Gaussian Quadrature rules give the exact approximation of integral (7).
@@
Theorem 2 shows the existence of good $\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ and their general form. In fact, it is also shown that Gaussian Quadrature is optimal: no other$\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ can achieve zero approximation error for higher degree polynomials $f(\lambda)$ [34]. However, it is often difficult to find these quadrature rules [36]. There are at least three questions in sequel:
@@
Now suppose that we have found the orthogonal polynomials $\left\{p_{k}(x)\right\}_{k=1}^{m}$ w.r.t. $\mu$. Recall that any orthogonal polynomial has the following "three-term" recursion [34]:
@@
[34] For the Gaussian Quadrature, $\{x_{j}\}_{j=1}^{m}$ are the eigenvalues of $J_{m}$ and $\{c_{j}\}_{j=1}^{m}$ are the squares of the first elements of the normalized eigenvectors of $J_{m}$.
The proof of Theorem 3 is based on Christoffel-Darboux relation [11]. Now, the remaining question is: how to find the Jacobian matrix $J_{m}$ of a sequence of orthogonal polynomials w.r.t. an unknown measure $\mu$? Note that we no longer need to answer question 1) if $J_{m}$ is found, since $J_{m}$ is sufficient for us to find the Gaussian quadrature rules. However, it seems impossible to find $J_{m}$ if no information of $\mu$ is provided. The good news is: when the $\mu$ is specified as in (8), there exists an efficient way to find $J_{m}$.
@@
When $\mu$ is specified as in (8), $J_{m}$ can be exactly found in $m$ steps using the Lanczos algorithm [47], as shown in Algorithm 4. This method takes a real symmetric matrix as input and returns a tridiagonal matrix. It was originally proposed to solve eigenvalue problems. Later, researchers found a deep connection between the Lanczos algorithm and orthogonal polynomials, which further connects this method to the Gaussian quadrature. The method (of finding the Gaussian quadrature by the Lanczos algorithm) is called the Lanczos quadrature [35, 6, 34]. An extremely elegant but highly nontrivial result is as follows:
@@
[34] Given a real symmetric matrix $H\in\mathbb{R}^{d\times d}$ and an arbitrary vector $v\in\mathbb{R}^{d}$ with unit Euclidean norm, we define the measure $\mu$ as in (8) based on this $H$ and $v$. Then $m$ steps of the Lanzcos algorithm return the Jacobian matrix $J_{m}$ of orthogonal polynomials w.r.t. to $\mu$.
@@
The full procedure of SLQ is shown in Algorithm 5. We note that SLQ is efficient in theory. Ubaru et al. [83] show that SLQ converges faster than any other polynomial expansion method for spectrum estimation (e.g., Chebyshev methods used in [2]). See [83, Theorem 4.1] for a formal statement.
We remark that there are at least four versions of the Lanczos algorithm in Step 3. Here, we adopt the version in Algorithm 4 since it is known to be the most numerically stable version [20, 74, 90]. Throughout this work, we choose $f\left(\cdot\right)$ as the Gaussian blurring function $f(\lambda):=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(t-\lambda)^{2}}{2\sigma^{2}}\right)$ for spectrum approximation. We plot the spectrum by sweeping $t$ from the minimal node to the maximal node in Gaussian Quadrature rules.
@@
Appendix D More Eperimental Details
D.1 Implementation Details on SLQ and Training Configurations